OpenStudy (anonymous):

Evaluate 3^2 + (7 - 2) • 4 - 6 over 3.

OpenStudy (anonymous):

7.6

OpenStudy (anonymous):

PEMDAS

OpenStudy (igreen):

\(\dfrac{3^2 + (7-2) • 4 - 6}{3}\) PEMDAS: P = Parenthesis E = Exponent M = Multiplication D = Division A = Addition S = Subtraction Parenthesis: \(\dfrac{3^2 + 5 • 4 - 6}{3}\) Exponent: \(\dfrac{9 + 5 • 4 - 6}{3}\) Multiplication: \(\dfrac{9 + 20 - 6}{3}\) Addition: \(\dfrac{29 - 6}{3}\) Subtraction: \(\dfrac{23}{3}\) So your final answer is: \(\dfrac{23}{3}\) or \(7 \! \! \dfrac{2}{3}\)

OpenStudy (anonymous):

same thing what i said just a fraction

OpenStudy (anonymous):

7.66 Repeating is correct. 23/3 or 7+2/3

OpenStudy (igreen):

@NaeNaeHoowaahh \(\hspace{80pt}\bbox[5pt, black,border:5px solid #00FF00]{\bf\huge\color{#00FF00}{\href{javascript:confirm('Your OpenStudy user has been issued a warning by iGreen. Please do not give out direct answers. Refer to our code of conduct at openstudy.com/code-of-conduct for more information.')}{Click~Here}}}\)

OpenStudy (anonymous):

ok sorry

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @medeline22